OpenStudy (anonymous):

Simplify -3x + 9 / -3 x − 3 x + 9 −x − 3 −3x − 3

OpenStudy (anonymous):

\[\frac{-3x + 9 }{ -3}\]

OpenStudy (anonymous):

factor the top out. Preferable you should factor out a 3 so you can cancel the 3 from the denominator

OpenStudy (anonymous):

\[\frac{-3x + 9}{-3} = \frac{-3(x - 3)}{-3} = \frac{\cancel{-3}(x - 3)}{\cancel{-3}} = x - 3\]

OpenStudy (anonymous):

Thank you guys =)

OpenStudy (anonymous):

there you go you have your answer.

OpenStudy (anonymous):

What you do here is factor out the -3 on top. Then, they cancel. You're welcome :)

OpenStudy (anonymous):

.....